# Social Media Isn’t Just Instagram: A Youth-Envisioned Platform for Meaningful Social Connections

### v1

@@ Abstract.
We conducted co-design workshops with 23 participants (ages 15–24) to explore how youth envision an ideal remote social connection. Using the Fictional Inquiry (FI) method within a Harry Potter-inspired narrative, we found that youth perceive a disconnect between platforms labeled as “social media” (like Instagram) and those where they actually experience meaningful connections (like Minecraft or Discord). Participants envisioned an immersive 3D platform that would bridge this gap by prioritizing meaningful social connections, enabling time that feels well spent through presence and immersion, natural individual expression, intuitive social navigation that leverages physical-world norms, and playful, low-stakes opportunities for gradual friendship development. We introduce the design framework of spatial integrity, which encompasses four dimensions of spatial affordances that facilitate meaningful social connections online. The FI method proved effective in generating innovative ideas while empowering youth by fostering a sense of hope and agency over the future of social media through their own design contributions.
@@ 1. Introduction
Most youth today use social media daily and spend significant time on platforms such as Instagram and TikTok (Anderson, 2023). However, these same youth often express deep dissatisfaction with how they spend this time—not because they want less social connection, but because they desire more meaningful online interactions (Kim et al., 2024c). They often feel constrained by the design patterns of archetypal social media platforms, which prioritize uninteresting and attention-grabbing content consumption over genuine connection (Landesman et al., 2024). This disconnect has contributed to broader societal concerns about social media’s impact on youth well-being (Oth, 2023).
Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023). This narrative has prompted restrictive policy responses worldwide: 41 U.S. states suing Meta (Lima-Strong and Nix, 2023), proposals for “warning labels” on social media use (Murthy, 2024), restrictive legislation like the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act” (Oth, [n. d.]), and outright bans on social media use for those under 16 in countries like Australia (Kim, 2024). The EU’s GDPR (gdp, [n. d.]) similarly limits teen access.
However, such alarmist narratives oversimplify a complex reality: the effectiveness of these narratives and policies (as well as the causal role of social media in youth mental health) remains unclear. Despite recognizing the issues raised in these narratives, the youth report more positive than negative personal experiences (Vogels, 2023), even as they acknowledge potential risks to others (Vogels, 2023). Rather than perpetuating fear-based narratives or attempting to completely rebuild existing platforms, researchers and teens themselves emphasize the importance of learning to navigate social media effectively (Wisniewski et al., 2012a; Wisniewski, 2018a; The Learning Network, 2025). Furthermore, such alarmist responses and outright restrictions may be misguided—they impede the development of resilience (Wisniewski, 2018b) and induce extreme and unhelpful anxiety and mistrust without fostering constructive actions (Kim et al., 2024a), ultimately leading to lower psychological well-being associated with social media use (Lee and Hancock, 2024).
This misalignment suggests that the solution lies not in the wholesale rejection of social media or attempts to limit youth access, but in expanding our understanding of what social platforms can be. Rather than trying to “fix” existing platforms or promote fear-based narratives, we need to broaden our vision of what social media could be. To explore alternative visions for social technologies that prioritize meaningful connection, we employed the Fictional Inquiry (FI) method (Dindler and Iversen, 2007) in co-design workshops with 23 participants (ages 15-24) experienced in 3D gaming environments (e.g., Minecraft (min, [n. d.]), Roblox (Roblox Corporation, [n. d.])). Participants were asked to envision a social media platform for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n. d.]). This magical context freed participants from current platform constraints while maintaining familiar social dynamics.
Our findings reveal two important patterns in how youth view and experience online social platforms. First, they make a clear distinction in their minds: platforms like Instagram are what they consider “social media”—where they spend significant time but rarely form meaningful connections. In contrast, they do not typically label platforms like Discord or Minecraft as “social media,” even though these are where they actually develop deep friendships and find their time meaningful. Second, because youth associate “social media” primarily with platforms like Instagram and its limitations, they become skeptical of social media’s overall potential to support authentic relationships. This narrow definition of social media, shaped by their experiences with Instagram-like platforms, leads them to dismiss the entire category as inherently flawed.
Participants identified several key elements missing from archetypal social media platforms: (1) emotional connection through presence and immersion, (2) diverse forms of individual expression that feel more natural, (3) intuitive social navigation—ranging from privacy to relationship development—that enables them to leverage norms from the physical world, and (4) playful, low-stakes opportunities to develop friendships. Although participants expressed deep dissatisfaction with the time they currently spend on archetypal social media platforms, they felt that spending the same amount of time on platforms designed to prioritize meaningful connections would feel valuable.
@@ 2. Related Work
2.1. Defining Social Media and its Role in Young People’s Lives
Social media does not yet have a widely agreed-upon definition. There is a range of interpretations, from broad (e.g., “any interactive communication medium that enables two-way interaction and feedback” (Kent, 2010)) to specific (e.g., “a different kind of Internet-based applications which build the ideological and technological foundations of Web 2.0, and allow users to create content and exchange it with other people through the Internet” (Kaplan and Haenlein, 2010)). While some platforms, such as Instagram (All, [n. d.]) or Facebook (fac, [n. d.]), are indisputably considered social media, others, such as YouTube (you, [n. d.]) and Discord (discord page, [n. d.]), are less archetypal. There are also different types of social media, and even such classifications have multiple interpretations (Kaplan and Haenlein, 2010; Sharma and Verma, 2018; Scott, 2015). Some major categories include social networking sites (i.e., “web-based services that allow individuals to (1) construct a public or semi-public profile within a bounded system, (2) articulate a list of other users with whom they share a connection, and (3) view and traverse their list of connections and those made by others within the system” (Boyd and Ellison, 2007)) and blogs (i.e., “a way to share that love with the world and encourage an active community of readers who comment on the posts of the author” (Scott, 2015)).
The range of definitions and classifications of social media reflects the wide array of user needs that social media addresses. From a uses and gratifications perspective (Katz et al., 1973), social media supports a variety of utilities, ranging from social interactions to information seeking and entertainment (Whiting and Williams, 2013). Even within such broad categories of uses and gratifications, there are diverse ways these needs are pursued. For example, within social interactions, there may be a focus on bridging social capital and weaker ties versus bonding social capital and stronger ties (Phua et al., 2017). Further, even on a single platform—Instagram, for instance—users engage in a range of activities, from photo sharing for self-promotion and disclosure (Menon, 2022) to more passive consumption of media for entertainment. In other words, Instagram may serve as a personal branding tool for some while functioning as a way to pass the time for others.
Despite the prevalent narrative that social media is harmful, the experiences of youth with these platforms are more nuanced. They face challenges such as social comparison (Yau and Reich, 2019), significant privacy concerns (Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014), pressures to curate their online presence (Yau and Reich, 2019), and feelings of meaningless time spent (Lukoff et al., 2018), along with more subtle discomforts (Landesman et al., 2024). However, these platforms also offer meaningful benefits. For instance, marginalized populations often derive substantial support from these spaces (Li et al., 2023; Bates et al., 2020), and social media serves as an important medium for identity development and self-expression (Davis, 2012; Lee et al., 2022). Given this complex relationship between youth and social media, researchers advocate for strategies that focus on building resilience and equipping teens to navigate these experiences effectively rather than solely aiming to restrict or mitigate harm (Wisniewski et al., 2012b; Wisniewski, 2018a; Agha et al., 2023).
@@ 2.2. Social Technology Design Explorations in HCI
HCI research has extensively explored designs that foster meaningful social interactions online, examining both platform-level innovations and individual features. Social media platforms such as BeReal (Kim et al., 2024c), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) have introduced unique affordances for sharing and connecting, with an emphasis on spontaneity, ephemerality, and creative self-expression. This line of research identified different ways to build and maintain relationships through features that prioritize authenticity and playful interaction. However, users have reported increasingly toxic interactions as they gain popularity, with design choices encouraging users to expand their networks and engage in compulsive behaviors, such as tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), which detract from their initial goals of fostering meaningful interactions.
@@
Beyond traditional social media, recent work has investigated spatial social technologies, which emphasize the value of immersive, embodied interactions. Social virtual reality (VR) platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide users with virtual spaces and embodiment for co-presence and shared experiences, facilitating interactions that feel more personal and tangible compared to 2D environments. Similarly, platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend elements of gaming and social networking, offering users the ability to navigate virtual environments, customize avatars, and participate in dynamic, context-rich group interactions. However, these platforms often are not generally recognized as social media in the traditional sense.
@@
What youth envision as ideal social media, both at the broader landscape level and the specific feature level remains unclear—especially when considering the possibilities offered by emerging technologies such as spatial computing platforms. Understanding these perspectives is crucial for reimagining the future of social media and ensuring it meets the evolving needs of its users.
2.3. Designing Beyond Doom Narratives
@@
Social media and design fixation. Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented challenge in design. While generating a wide range of divergent ideas is a critical first step in innovative design, we often impose unnecessary restrictions on our thinking. This self-limiting cognitive process often confines the design process to address localized problems rather than exploring broader, more optimal solutions.
@@
Social media design is no exception. Given the ad-based revenue model, platforms are incentivized to increase time spent on their apps. However, instead of achieving this through meaningful utility and engagement that genuinely enhances users’ experiences, they often resort to less healthy, compulsive strategies, such as friend recommendations (Kim et al., 2024c) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)). These tactics drive superficial interactions rather than fostering deeper connections, leaving users dissatisfied and reinforcing the perception that all social media designs and user experiences are fundamentally the same (Sundaram, [n. d.]; Pardes, 2020).
@@
The Fictional Inquiry (FI) method. FI is an exploratory co-design method that enables participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007). At its core, FI creates an environment where people can imagine beyond present limitations by engaging with carefully crafted scenarios. These scenarios typically incorporate immersive elements such as physical artifacts or role-playing activities, with participants taking on generative roles that help them envision futuristic or fictional settings. The method is particularly effective at helping participants overcome design fixation by encouraging them to draw analogies from fictional scenarios and imagine freely outside real-world constraints. While FI shares the exploratory nature of other ideation techniques, it is distinguished by its emphasis on developing futuristic design materials through collaborative dialogue between designers and users (Dindler and Iversen, 2007).
@@ 3. Method
We conducted co-design workshops with 23 participants between the ages of 15 and 24 to examine how we might envision an ideal social media platform free from real-world constraints and from design fixation on archetypal social media platforms. All participants had prior experience with 3D gaming and familiarity with Harry Potter (Rowling, 2015; har, [n. d.]), a popular series of young adult fantasy novels (and movies based on these stories) in which a subset of the population possesses magical powers. During the workshop, we used the FI method to scaffold participants in generating design ideas.

### v2

# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

@@ Abstract.
We investigate youth visions for ideal remote social interactions, drawing on co-design interviews with 23 participants (aged 15-24) experienced with 3D gaming environments. Using a Fictional Inquiry (FI) method set in the Harry Potter universe, this research reveals that young people desire social media that functions more like immersive, navigable shared social spaces. Across these interviews, participants identified six key priorities for meaningful social connection over social media: intuitive social navigation, shared collaborative experiences, communal environments fostering close relationships, flexible self-presentation, intentional engagement, and playful social mechanics. We introduce the spatial integrity framework, a set of four interrelated design principles: spatial presence, spatial composition, spatial configuration, and spatial depth. Together, these principles outline how online spaces can be designed to feel more like meaningful environments, spaces where relationships can grow through shared presence, movement, and intentional interaction. Participants also described the FI process itself as meaningful, not only for generating new ideas but for empowering them to imagine and shape the future of social media.
@@ 1. Introduction
Meaningful social connections play a critical role in young people’s development and well-being (Hawkley and Cacioppo, 2010), and many young people turn to social media platforms to build and maintain these relationships (Allen, 2019; Bohn, [n. d.]). However, recent trends indicate that social media platforms are increasingly becoming less focused on facilitating personal interactions (Nielsen and Fletcher, 2023; Chayka, 2025). Initially designed around sharing personal updates with friends, these platforms now often emphasize entertainment, news, and engaging larger, less intimate audiences. Meta’s own data illustrates this shift: between 2023 and 2025, the share of time users spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley et al., 2025). Mark Zuckerberg described this change as aligned with the company’s focus on supporting users “the general idea of entertainment and learning about the world and discovering what’s going on” (Chayka, 2025).
This shift away from personal interactions raises concerns about the future of meaningful online relationships, particularly for younger users who have grown up with social media as a central part of their social lives. Most youth today engage with social media daily, spending substantial time on platforms like Instagram and TikTok (Anderson, 2023). Despite this high engagement, many young users report dissatisfaction with the quality of their social interactions, not because they wish to reduce their usage, but because they expect these platforms to better support their social connections (Kim et al., 2024b). Current platform designs frequently prioritize passive consumption of content over active interaction with close friends (Landesman et al., 2024), contributing to broader concerns about social media’s impact on youth well-being (Oth, 2023).
Given the central importance of social connections, particularly for youth, and recognizing social media’s vital role in facilitating these interactions, it is increasingly important to explore new design possibilities. Although young people rely heavily on social media, they frequently express dissatisfaction with their current experiences, indicating a significant gap between their social needs and the interactions enabled by existing platforms. Current mainstream social media platforms employ highly standardized design patterns, which limit our collective imagination and contribute to feelings of pessimism among both younger and older users. Consequently, there is a compelling need for innovative designs that complement existing platforms by placing a stronger emphasis on meaningful interpersonal connections. This prompts the critical question of understanding what types of social interactions current designs might be overlooking or inadequately supporting.
To address this question, we conducted a series of co-design interviews using the Fictional Inquiry (FI) method (Dindler and Iversen, 2007). We involved 23 participants aged 15–24 who have experience with 3D gaming environments like Minecraft (min, [n. d.]) and Roblox (Roblox Corporation, [n. d.]). Participants were asked to collaboratively imagine a social media platform specifically designed for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n. d.]). This imaginative context allowed participants to step away from the constraints of existing platforms and envision more varied and meaningful social interactions. Additionally, because 3D gaming environments are widely recognized as effective spaces for youth socialization alongside social media (Gentile et al., 2009; Granic et al., 2014; Lenhart, 2015c), we anticipated that participants could leverage insights from their gaming experiences to propose novel social media interactions. Specifically, our research aims to answer the following question: What types of interactions do youth familiar with 3D gaming environments envision as ideal for supporting meaningful remote social connections?
Our findings reveal that youth with 3D gaming experiences envision ideal social media platforms not as immersive, navigable environments designed to feel more like neighborhoods, community centers, and shared social spaces. Participants described six priorities: (1) intuitive social navigation modeled on real-world movement and presence; (2) shared experiences through collaborative activities and casual co-presence; (3) communal environments that prioritize close relationships, conversations, and low-pressure interaction; (4) flexible self-presentation through avatars, personal spaces, and contextual identity management; (5) intentional engagement that replaces passive scrolling with purposeful navigation; and (6) playful, low-stakes social mechanics inspired by gaming. Together, these visions highlight the youth’s desire for social media that feels more relational, meaningful, and personally rewarding.
@@ 2. Related Work
2.1. Meaningful Social Connection in Online Environments
2.1.1. Defining Meaningful Social Interaction
Litt et al. (Litt et al., 2020) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” While these interactions are described as “deeply subjective,” the primary distinction from non-meaningful ones lies in the presence of impact. This impact extends “beyond the immediate interaction itself,” typically involving positive emotions, the exchange of information, or concrete benefits.
The study identified three key types of impact: “emotional impact” (involving feelings, mood changes, or relationship strengthening), “informational impact” (relating to the exchange of knowledge, advice, or understanding), and “tangible impact” (providing concrete benefits such as goods or services). In contrast, non-meaningful interactions were often explicitly described as lacking any such impact.
Factors like the presence of strong ties (e.g., friends and family) and community ties, as well as participation in shared activities, were found to be associated with meaningfulness. However, these appeared to act more as facilitators that enable impact rather than as the primary source of meaning. Other attributes quantitatively linked to greater meaningfulness included whether interactions were planned in advance, memorialized (for example, through photos or videos), and synchronous. Interestingly, the study found that the communication medium itself, whether online or offline, was less important in predicting meaningfulness after accounting for these other characteristics.
While in this paper we treat meaningful social interaction as any interaction that participants say they would actively seek out in an ideal scenario. reflecting their subjective or perceived meaningfulness, we expected these themes to echo in how our participants talked about their experiences.
2.1.2. Social Connection as a Key User Need Left Unmet by Today’s Platforms
Social media platforms are often seen as a primary means of connecting with others, especially for young people (Boyd, 2014a). However, despite the widespread use of these platforms, many users report feeling lonely or disconnected (Twenge et al., 2019; Primack et al., 2017). This disconnect highlights a significant gap between the intended purpose of social media and the actual experiences of users.
Social connection is just one of the many reasons people turn to social media, but it remains one of the most important. From staying in touch with close friends to meeting new people or maintaining community ties, users often come to these platforms seeking connection. Mainstream platforms like Instagram (All, [n. d.]), YouTube (you, [n. d.]), and Discord (discord page, [n. d.]) support some of these needs, but they are designed to address a wide range of user motivations, including entertainment, information seeking, and self-promotion (Whiting and Williams, 2013; Menon, 2022). From a uses and gratifications perspective (Katz et al., 1973), even within a single platform, users pursue diverse goals, ranging from building social capital (Phua et al., 2017) to engaging in general media consumption (Menon, 2022). As such, although people may approach these platforms with the intention of connecting socially, they may encounter experiences that prioritize other forms of engagement, potentially leaving their social needs unmet.
This gap highlights the importance of developing a clearer understanding of what it means to support meaningful social connection in online environments. While social media offers both risks, such as social comparison and privacy concerns (Yau and Reich, 2019; Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014b), and benefits, such as support for marginalized groups (Li et al., 2023; Bates et al., 2020; Steeds et al., 2025) and opportunities for identity development and self-expression (Davis, 2012; Lee et al., 2022), social connection remains a central motivation for many users. Particularly for young people, who increasingly rely on digital platforms for social interaction, it is critical to move beyond treating connection as one of many competing platform goals and toward a more focused understanding of how platforms can meaningfully support users’ social relationships.
@@ 2.2. Social Technology Design Explorations in HCI
2.2.1. Designing for Meaningful Social Connections
HCI research has extensively explored designs that foster meaningful social interactions online, examining both platform-level innovations and individual features. Social media platforms such as BeReal (Kim et al., 2024b), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) have introduced unique affordances for sharing and connecting, with an emphasis on spontaneity, ephemerality, and creative self-expression. This line of research identified different ways to build and maintain relationships through features that prioritize authenticity and playful interaction. However, users have reported increasingly toxic interactions as they gain popularity, with design choices encouraging users to expand their networks and engage in compulsive behaviors, such as tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), which detract from their initial goals of fostering meaningful interactions.
@@ 2.2.1. Designing for Meaningful Social Connections
Beyond archetypal social media, recent work has investigated spatial social technologies, which emphasize the value of immersive, embodied interactions. Social virtual reality (VR) platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide users with virtual spaces and embodiment for co-presence and shared experiences, facilitating interactions that feel more personal and tangible compared to 2D environments. Similarly, platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend elements of gaming and social networking, offering users the ability to navigate virtual environments, customize avatars, and participate in dynamic, context-rich group interactions. However, these platforms often are not generally recognized as social media in the generic sense.
@@
What youth envision as ideal social media, both at the broader landscape level and the specific feature level remains unclear, especially when considering the possibilities offered by emerging technologies such as spatial computing platforms. Understanding these perspectives is crucial for reimagining the future of social media and ensuring it meets the evolving needs of its users.
2.2.2. Video Games as a Source of Meaningful Social Connection
Video games are increasingly recognized as valuable environments for fostering meaningful social connection, particularly among youth. Large-scale survey research shows that video games serve as a major venue for the development and maintenance of friendships, especially for boys (Lenhart, 2015c). Teens not only play games with friends they know in person but also frequently make new friends through online gameplay (Lenhart, 2015a), with more than half of teens reporting they have met at least one friend online (Lenhart, 2015a, b). These interactions extend beyond gameplay itself, supporting ongoing social relationships through shared experiences, communication, and collaboration (Lenhart, 2015c).
Research also highlights the social and emotional benefits of gaming, including opportunities for cooperation, empathy, and prosocial behavior (Gentile et al., 2009; Granic et al., 2014). Certain game designs promote prosocial outcomes by encouraging players to help and support others, both in-game and beyond (Gentile et al., 2009). Moreover, games are not limited to surface-level entertainment; they can also offer deeper, eudaimonic experiences that involve meaning-making, personal growth, and feelings of connection with others (Daneels et al., 2023). These findings suggest that video games can be powerful tools for fostering meaningful social connections, particularly among youth, and highlight the potential for game-like features to enhance social media experiences.
2.3. Recognizing and Overcoming Social Media Design Fixation
@@
Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented challenge in design. While generating a wide range of divergent ideas is a critical first step in innovative design, we often impose unnecessary restrictions on our thinking. This self-limiting cognitive process often confines the design process to address localized problems rather than exploring broader, more optimal solutions.
@@
Social media design is no exception (Sundaram, [n. d.]; Pardes, 2020; Pitt, 2021; Nielsen and Fletcher, 2023; VanDyke, 2020). Given the ad-based revenue model, platforms are incentivized to increase time spent on their apps (Oth, [n. d.]). However, instead of achieving this through meaningful utility and engagement that genuinely enhances users’ experiences, they often resort to less healthy, compulsive strategies, such as friend recommendations (Kim et al., 2024b) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)). These tactics drive superficial interactions rather than fostering deeper connections, leaving users dissatisfied and reinforcing the perception that all social media designs and user experiences are fundamentally the same (Sundaram, [n. d.]; Pardes, 2020).
2.3.1. Fictional Inquiry (FI) as a Tool for Imaginative Design Exploration
@@
FI is an exploratory co-design method that enables participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007). At its core, FI creates an environment where people can imagine beyond present limitations by engaging with carefully crafted scenarios. These scenarios typically incorporate immersive elements such as physical artifacts or role-playing activities, with participants taking on generative roles that help them envision futuristic or fictional settings. The method is particularly effective at helping participants overcome design fixation by encouraging them to draw analogies from fictional scenarios and imagine freely outside real-world constraints. While FI shares the exploratory nature of other ideation techniques, it is distinguished by its emphasis on developing futuristic design materials through collaborative dialogue between designers and users (Dindler and Iversen, 2007).
@@
Further, one of the most powerful ways to overcome design fixation is analogical transfer (Novick, 1988; Buchanan, 1992). We see video games as a rich source of analogies for social media design, offering unique affordances and interaction patterns that can inspire new approaches to fostering meaningful social connections. By examining how video games facilitate social interactions, we can identify design patterns that can be adapted to social media contexts, helping to break free from design fixation and explore novel alternatives.
@@ 3. Method
We conducted co-design interviews with 23 participants between the ages of 15 and 24 to examine how we might envision an ideal social media platform free from real-world constraints and from design fixation on archetypal social media platforms. All participants had prior experience with 3D gaming and familiarity with Harry Potter (Rowling, 2015; har, [n. d.]), a popular series of young adult fantasy novels (and movies based on these stories) in which a subset of the population possesses magical powers. During the interviews, we used the FI method to scaffold participants in generating design ideas.